OpenStudy (anonymous):

Find the first of the nth roots of the complex number. Leave answer in polar form and use degrees for theta. 81 n=4 ________________ I got to this part. x^4=81 theta=0,90,180, 270

OpenStudy (dumbcow):

is the complex number 81?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i must find 4 roots of 81

OpenStudy (dumbcow):

polar form of 81: \[81(\cos 0 + i \sin0)\] 4th roots are \[\sqrt[4]{81}(\cos \frac{0+360k}{4} + i \sin \frac{0+360k}{4})\] where k = 0,1,2,3